OpenStudy (thunder33):

The vertex form of the equation of a parabola is y = (x - 5)2 + 16. What is the standard form of the equation?

OpenStudy (welshfella):

first expand (x - 5)^2 (x - 5)( x - 5) = ?

OpenStudy (welshfella):

first multiply the x - 5 in second parentheses by x then multiply the x - 5 by - 5

OpenStudy (welshfella):

x(x - 5) - 5(x - 5) + 16 = x^2 - 5x -5x .. can you continue?

OpenStudy (thunder33):

i got y=x^2-10x+41

OpenStudy (welshfella):

correct